OpenStudy (anonymous):

Taylor is making a cookie cake that has a diameter of 13 in. What is the area of one-half of the cookie cake? 66.37 265.46 65.73 48.4

OpenStudy (anonymous):

Try: http://www.mathopenref.com/circlearea.html You know the diameter, but you could divide by 2 to get the radius. Then use the formula. You will need a calculator to get the answer since they have given approximations to the area.

OpenStudy (anonymous):

\[\frac{1}{2} \pi \left(\frac{13}{2}\right)^2=\frac{169 \pi }{8}=66.37 \]

OpenStudy (anonymous):

first find the area of your circle using the formula \[\pi*r^2\] then multiply by 1/2 to find the area of half the circle